how do i solve 1/3 x+ 3/4x=13
and
-2x + 4 > 12

Answer by stanbon(75887)

solve 1/3 x+ 3/4x=13
Multiply both sides by 12 to get:
4x + 9x = 13*12
13x = 13*12
x = 12
-------------------------
and
-2x + 4 > 12
Add 2x to both sides to get:
4 > 2x+12
Divide thru by 2 to get:
2 > x+6
Subtract 6 from both sides to get:
x < -4
==============
Cheers,
Stan H.